OpenStudy (anonymous):

I need help on how to work out this simultaneous equation: 3L+25=19.50 4L+45=30

OpenStudy (anonymous):

I've already done 7L=49.50 L=7.07 not so sure what to do next?

OpenStudy (mathmale):

Hope you won't mind, but I'd prefer you demo how you got 7L=49.50. That does not look appropriate to me. You're working towards solving a system of linear equations. There are three relatively simple methods to choose from: (1) elimination by addition/subtraction; (2) elimination by substitution; and (3) graphing. Which method would you like to use first?

OpenStudy (anonymous):

Well i looked up on BBC bitesize how to do it and i don't think I've understood properly, i done 3L+4L=7L then i done 19.50+30=49.50 and then i got 7L=49.50

OpenStudy (mathmale):

I'm afraid I have to disagree with that. 3L+25=19.50 becomes 3L = 19.50 - 25.00, and 4L+45=30 becomes 4L = 30 - 45 Add the left sides of these two equations together. Then add the right sides together. Are you certain that you've copied the problems down correctly? I'd think there was another variable, such as W, involved. Note that if you have copied the problem correctly, 4L = 30 - 45 = -15, so L =-15/4.

OpenStudy (mathmale):

Again, I ask you to check that you've copied the problem down correctly. If you have, please choose one of the three methods of solution that I've proposed.

OpenStudy (anonymous):

So 7L+70= What do i put for the equals?

OpenStudy (mathmale):

Again, I ask you to check that you've copied the problem down correctly. (Please.) I suspect that there's another variable in this system of equations.